Find two complex numbers that have a sum or -6 and a product of 90
Answer by khwang(438)

Let these two numbers be the roots of a quadratic equation.
Then the equation should be x^2 + 6x +90 = 0.
By quadratic formula, we have x = -3 + sqrt(9-90) = -3 + 9i
or -3 - 9i.
Thus, we get these two required complex numbers.
Kenny